Consider a car moving along a straight horizontal road with a speed of 72 km/h. If the coefficient of static friction between the tyres and the road is 0.5, the shortest distance in which the car can be stopped is (taking g = 10 m/s2)ย 


(a)ย ย ย  30m


(b)ย ย ย  40m


ย ย ย ย ย ย ย  (c)ย ย ย  12m


(d)ย ย ย  20m


Grade:11

1 Answers

Navjyot Kalra
10 years ago

(b)
Here u = 72 km/h = 20 m/s; v = 0;
a =- ?g = -0.5 × 10 =-5m/s2
As v2 = u1 + 2as,
? s=  ((v^2- u^2 ))/2a  =((0- (20)^2)/(2× (-5) )=40 m